b.) one mole of photons of 3.91 x 1019 Hz light?

what is the E associated with

b.) one mole of photons of 3.91 x 1019 Hz light?
Answer #1

Given

Frequency = 3.91 \times 1019 Hz

we know

h = Plancks constant = 6.626 \times 10-34 J.s

we have formula

Energy of photon = h \times frequency

= 6.626 \times 10-34 J.s \times 3.91 \times 1019 Hz

= 2.59 \times 10-14 J

1 mole = Avogadros number of photons = 6.022 \times 1023 photons

so the energy of one mole of photon is

Energy = 2.59 \times 10-14 J \times 6.022 \times 1023 mol-1  

= 1.56 \times 1010 J/mol

= 1.56 \times 107 kJ/mol